OpenStudy (anonymous):

write the equation in standard form that passes through (-5,1) and (6,-2)

OpenStudy (anonymous):

well standard form is Ax + By = C

OpenStudy (anonymous):

but you must first get the slope intercept form which is y = mx + b

OpenStudy (anonymous):

then find the y intercept

OpenStudy (anonymous):

by plugging in any of the two points

OpenStudy (anonymous):

ya find gradient 1st before anythg else

OpenStudy (anonymous):

3/5 is the slope

OpenStudy (anonymous):

so you have y = 3/5x + b so if you plug in the point (-5,1) you'll get the equation 1 = 3/5(-5) + b 1 = -3 + b b = 4 so y = 3/5x + 4 would be the slope intercept form

OpenStudy (anonymous):

then you can just change it into standard form

OpenStudy (anonymous):

you can change this form into standard form by moving the 3/5x to the other side

OpenStudy (anonymous):

so you'll get -3/5x + y = 4